# Unsupervised Representation Learning for 3D Mesh Parameterization with Semantic and Visibility Objectives

### v2

@@ A.3 Additional Experiments and Results
Figure 14: Ablation studies on the proposed partitioning strategy (ShDF vs. SAMesh vs. PartField) of simple meshes. We replace our default Shape Diameter Function (ShDF) partitioner Shapira et al. (2008) with SAMesh Tang et al. (2024) and PartField Liu et al. (2025) while keeping the remainder of the partition-and-parameterize pipeline identical. Qualitatively all three partitioners can produce part decompositions that align with geometric structure across simple meshes, resulting compact, low-distortion per-part UV charts. However, it seems that PartField produces more accurate and more detailed semantic segmentations than ShDF and SAMesh in some cases (e.g., the feet and the tail are correctly detected, whereas they are not detected properly in SAMesh).
@@
Figure 15: Ablation studies on the proposed partitioning strategy (ShDF vs. SAMesh vs. PartField) of complex meshes with high number of triangle faces. We replace our default Shape Diameter Function (ShDF) partitioner Shapira et al. (2008) with SAMesh Tang et al. (2024) and PartField Liu et al. (2025) while keeping the remainder of the partition-and-parameterize pipeline identical. Qualitatively all partitioners produce part decompositions that align with geometric structure across complex models. However, PartField appears to produce more accurate and more compact semantic segmentations than ShDF and SAMesh in some cases. For example, PartField correctly groups the entire set of fan blades into a single segmented part, whereas SAMesh and ShDF split each blade into a separate part. This is not completely wrong, but it results in too many small charts after per-part UV parameterization, which may be undesirable for some downstream tasks.
@@
Figure 16: Examples of four of the eleven questions in our user‐study questionnaire. To evaluate semantic- and visibility-awareness of the proposed method, we conducted a user-preference study with 115 participants. We distributed the questionnaire to two groups: experts and general participants. Of the 115 participants, 70 are general participants (including graduate students with computer science and engineering backgrounds) and 45 are experts: 31 software engineers, 3 project managers, 2 product owners, 5 UV/layout artists, and 4 modeling artists, all working in the Media and Entertainment industry for film and games. Each participant completed 11 comparisons between textured 3D shapes and UV parameterizations produced by our method, FlexPara Zhao et al. (2025), OptCuts Li et al. (2018), Autodesk Maya Autodesk, Inc. (2025), Blender Blender Foundation (2025), and xatlas jpcy (2025). For each comparison, participants were rated each result according to three visual criteria: (i) texture-pattern smoothness/continuity, (ii) semantic alignment (how well colors correspond to meaningful parts), and (iii) seam visibility (how well seams are placed in occluded/less-obvious regions). The semantic-aware and visibility-aware evaluations comprised 5 and 6 questions, respectively, for a total of 11 comparisons per participant. Tables 2 and  10 report the percentage of expert and general participant preferences for each method, respectively. As the tables show, our proposed method is strongly preferred over the baselines by both expert industry-level users and general users.

### v3

@@ A.3 Additional Experiments and Results
Figure 13: Ablation studies on the proposed partitioning strategy (ShDF vs. SAMesh vs. PartField) of simple meshes. We replace our default Shape Diameter Function (ShDF) partitioner (Shapira et al., 2008) with SAMesh (Tang et al., 2024) and PartField (Liu et al., 2025) while keeping the remainder of the partition-and-parameterize pipeline identical. Qualitatively all three partitioners can produce part decompositions that align with geometric structure across simple meshes, resulting compact, low-distortion per-part UV charts. However, it seems that PartField produces more accurate and more detailed semantic segmentations than ShDF and SAMesh in some cases (e.g., the feet and the tail are correctly detected, whereas they are not detected properly in SAMesh).
@@
Figure 14: Ablation studies on the proposed partitioning strategy (ShDF vs. SAMesh vs. PartField) of complex meshes with high number of triangle faces. We replace our default Shape Diameter Function (ShDF) partitioner (Shapira et al., 2008) with SAMesh (Tang et al., 2024) and PartField (Liu et al., 2025) while keeping the remainder of the partition-and-parameterize pipeline identical. Qualitatively all partitioners produce part decompositions that align with geometric structure across complex models. However, PartField appears to produce more accurate and more compact semantic segmentations than ShDF and SAMesh in some cases. For example, PartField correctly groups the entire set of fan blades into a single segmented part, whereas SAMesh and ShDF split each blade into a separate part. This is not completely wrong, but it results in too many small charts after per-part UV parameterization, which may be undesirable for some downstream tasks.
@@
Figure 15: Examples of four of the eleven questions in our user‐study questionnaire. To evaluate semantic- and visibility-awareness of the proposed method, we conducted a user-preference study with 115 participants. We distributed the questionnaire to two groups: experts and general participants. Of the 115 participants, 70 are general participants (including graduate students with computer science and engineering backgrounds) and 45 are experts: 31 software engineers, 3 project managers, 2 product owners, 5 UV/layout artists, and 4 modeling artists, all working in the Media and Entertainment industry for film and games. Each participant completed 11 comparisons between textured 3D shapes and UV parameterizations produced by our method, FlexPara (Zhao et al., 2025), OptCuts (Li et al., 2018), Autodesk Maya (Autodesk, Inc., 2025), Blender (Blender Foundation, 2025), and xatlas (jpcy, 2025). For each comparison, participants were rated each result according to three visual criteria: (i) texture-pattern smoothness/continuity, (ii) semantic alignment (how well colors correspond to meaningful parts), and (iii) seam visibility (how well seams are placed in occluded/less-obvious regions). The semantic-aware and visibility-aware evaluations comprised 5 and 6 questions, respectively, for a total of 11 comparisons per participant. Tables 2 and  10 report the percentage of expert and general participant preferences for each method, respectively. As the tables show, our proposed method is strongly preferred over the baselines by both expert industry-level users and general users.